OpenStudy (anonymous):

what are the solutions to: |x−1|+|x−2|+|x−3|=4 ?

OpenStudy (perl):

hmmm

OpenStudy (perl):

if x>=3 , then we have (x-1) + (x-2) + (x-3) = 4 3x - 6 = 4 3x = 10 x = 10/3 . So that is one solution

OpenStudy (perl):

if 2<= x < 3 we have (x-1) + (x-2) + -(x-3) = 4

OpenStudy (perl):